OpenStudy (anonymous):

what is the point-slope equation of the line with a slope of -1/3 line that goes through the point (4,-5) answers y-5=1/3(x+4) y+5=1/3(x-4) y+5=-1/3(x-4) y-5=-1/3(x+4) We can eliminate A and B already since the slope is -1/3. To check to see which works do I do a test point with (0,0)?

OpenStudy (misty1212):

no just write \[y-y_1=m(x-x_1)\] which for you is \[y-(-5)=-\frac{1}{3}(x-4)\]

OpenStudy (misty1212):

the only tiny sophistication here is that \((--(-5)=+5\)

OpenStudy (misty1212):

actually i meant \(-(-5)=+5\) go with C

OpenStudy (anonymous):

OK thank you very much @misty1212

OpenStudy (misty1212):

you are welcome :)